Shift Summary for Owl shift, 29/Sep/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Wed, 29 Sep 1999 07:14:15 -0500

Shift Summary for Owl shift, 29/Sep/99

On Shift: P. Shanahan, M. Barrio

Hours of beam available at all intensities: 6.0
Hours of beam available at nominal intensity: 6.0
Hours of E799 data (special runs not included): 5.9
Hours of special runs (laser scans, muon runs...): 0.1

Accounting of lost time:
Beams was down at the beginning of the shift. Got beam consistenly
after 1:00, aproximately. Had some kfrending to do which added some
extra dead time.

========================================================================
Runs ended this shift:
Run 14754 - 17.3 M events - BEAM_799 - run didn't stop nicely,
changed tapes to avoid EOF
marker problems.
Run 14756 - ~200k events - BEAM_MUON - muon run.

Tapes brought to FCC this shift:
none.

========================================================================
Accesses during shift:
none.

========================================================================
Overview of shift:
Muon run; E799 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14754 0:00 No beam as of 23:28. Tried to stop this run
but the run control hung up.
Dart-fresh-started and changed tapes
because we don't know if the EOF mark was
written. New tapes: UP*046-048; UPS016
Tried to start a cosmic run but...

0:40 MCR calls, we'll have beam soon.

Changed laser delay from 19ns to 17ns in ktrig.

14756 0:50 Beam comes back. We go to muon mode for a muon
run (what else?).

Checked laser timing after correction, it's fine
now, but mean laser TDC (in lsr#time) is now
-3.4ns, which isn't consistent with TDC values
of several weeks ago -> RF timing shift might
be the culprit.

14756 1:00 Finished muon run (run # 14756) with circa 200k
events.

14757 1:10 Initialized E799 run. CHANGE OF POLARITY (neg ->
pos).

1:20 Start of run. Missing FFFF word errors in crate
53 (no wonder). Had to fix_disp a coupla times
too.

1:30 SDAQ complains that check_scaler is not running.
Indeed it isn't running, there are some typos
in the map file (the var names were not right).

2:00 Some typos in the namelist that defines the
"detector histos" which made a donkey appear
instead of the TRD trigger plot. Again Mista
Shanahan comes to the rescue and fixes the nml.

2:45 V1-East source seems to have a problem, it's not
counting, but other sources are (which indicates
it's not a detector problem but something in the
OR-ing for this source). However, the trigger
defs only have a 2V condition, so this source
is not being used. See logbook 21, page 18 for
more details.

4:40 Peter gets the check_scaler tcl to work.
We get constant B13, B15 and RC10 errors.
RC10 timing seems to be drifting around,
BobT and Elizabeth know about this.

5:40 Call MCR: SEM too high (~.5E12) for some 30
minutes now, Max will fix the thang; Asked about
drifting Acc90 rates, he'll look into it too.

Missing-FFFF-kfrend-ing galore.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift